pick a combination of one of the following molecules, one mass quantity, and one total solution volume from this table.

Molecule

Mass

Volume

NaCl

2 g

150 ml

CaBr2

6 g

3 L

Li3N

1 g

500 ml

CaO

7 g

14 L

For example, one possible combination would be 2 grams of NaCl in 500 ml of solution. Then, give each of the following for your unique combination being sure to show all of your work:

  1. undefined%mass/volume concentration
  2. undefinedMolarity
  3. undefinedOsmolarity
Answer #1

Consider 2g of NaCl in 500ml of solution.

1. Concentration = Mass/Volume = 2g/500ml = 0.004 g/ml

2. Molarity = No of moles of solute in 1 L solution.

Here, mol wt of NaCl is 23(for Na) + 35.5(for CL) = 58.5g

Now moles of solute = Given wt/Mol wt = 2/58.5 = 0.0342 moles.

So Molarity = No of moles of NaCl/ vol of solution in litres

= 0.0342/0.5 = 0.0684 M

3. Osmolarity = Molarity * no of active ions in 1 mol of solution = 0.0684 * 2 (since in 1 mole 2 ions would be present i.e., Na+ and Cl- ) = 0.1368 Osmol/L